1. A clothes dryer at home has a power of 4, 500 watts and runs on a special 220 V household circuit. a) What is the current through the dryer? b) How many kilowatt-hours of electricity are used by the dryer if it is used for 4 hours in one week? 



1
Expert's answer
2022-05-07T14:40:36-0400

(a)

"I=\frac{P}{V}=\frac{4500}{220}=20\:\rm A"

(b)

"W=Pt\\
=4.5\:\rm kW\times 4\: hours\times 7=126\: kW\cdot h"